# RISK: A Framework for GUI Agents in E-commerce Risk Management

### v1

@@ 6 Conclusion
In this work, we aim to address the critical challenge of automating e-commerce risk management tasks that involve dynamic, multi-step web interactions. Specifically, we propose the RISK framework, which incorporates a domain-specific dataset RISK-Data, a benchmark RISK-Bench, and a novel RFT approach RISK-R1 that comprises a comprehensive reward function and optimization objective to guide the model’s learning process. The experimental results confirm that RISK-R1 outperforms existing methods, showing a 6.8% improvement in single-step and an 8.8% improvement in multi-step performance, as well as achieving a top task success rate of 70.5% in real-world web environments. Our work provides a scalable, domain-specific solution for automating complex web interactions in high-stakes compliance and risk management tasks.
@@ 7 Ethics Statement
In this work, we ensure ethical compliance by sourcing all data exclusively from publicly available websites, with no personally identifiable information (PII) or sensitive data included. Strict data anonymization protocols are implemented to safeguard user privacy and address potential concerns.
Appendix A Appendix
A.1 Limitations and Future Work
Limitations. Although RISK-R1 demonstrates superior performance in e-commerce risk management tasks, there are still some limitations. (1) The current multi-step trajectories in RISK-Data are mainly used in SFT, while RFT only utilizes single-step trajectories due to GPU memory constraints. This may limit the model’s ability to fully learn multi-step decision-making processes. (2) Although we have incorporated process reweighting in our RFT framework to simulate an offline multi-step webpage interaction process, it may not fully capture the complexities and diversity of real-world scenarios. An online reinforcement learning framework could be more effective in this regard.
Future Work. Given the limitations mentioned above, we plan to address them in future work. We note that Shi et al. (2025) proposes a mobile GUI agent framework that leverages reinforcement learning in an online environment. We intend to adapt this framework to web-based GUI agents, enabling the model to learn directly from real-time interactions. Through this approach, GPU memory constraints can be alleviated and the model’s multi-step decision-making capabilities can be further enhanced. Additionally, we plan to build a high-concurrency cluster of browser environments to collect more diverse and complex multi-step instances, further enriching RISK-Data.
A.2 Task Definition
@@
A.2.1 External Information Retrieval and Extraction for Risk Intelligence
@@
A.2.2 Website Authenticity Verification for Risk Intelligence
A.3 Multi-step Trajectory Statistics
@@ A.4 Action Definition
Table 7: All actions and their definitions used in RISK (Part 2).
Action Definition
@@
A.5 Visualization of Weight Curve for Process Reweight
@@ A.6 Experimental Details
Evaluation Metrics. In offline single-step trajectory evaluations, we use the accuracy of tool calls as the evaluation metric, where a tool call is considered correct if its F1 score with the ground truth tool call exceeds 0.5. In offline multi-step trajectory evaluations, we use the task success rate as the evaluation metric, where a trajectory is considered successfully completed if all tool calls in the trajectory are correct. In online evaluations, we use the task completion rate and task success rate as the evaluation metrics, where the task completion rate is the percentage of tasks completed within a limited number of steps (set to 20), and the task success rate is the percentage of tasks successfully completed.
A.7 Ablation Study

### v2

@@ 6 Conclusion
This work tackles the automation of e-commerce risk management tasks that require dynamic, multi-step web interactions. We propose the RISK framework, including a domain-specific dataset (RISK-Data), a benchmark (RISK-Bench), and a new RFT method (RISK-R1) with a comprehensive reward function and optimization objective to guide learning. Experiments show that RISK-R1 offers a scalable, domain-specific solution for web interactions in high-stakes compliance and risk management.
7 Limitations
Although RISK-R1 demonstrates strong performance in e-commerce risk management tasks, several limitations remain: (1) Limited utilization of multi-step trajectories. In RISK-Data, multi-step trajectories are mainly used for supervised fine-tuning (SFT). By contrast, due to GPU memory constraints, reinforcement fine-tuning (RFT) is currently conducted primarily with single-step trajectories. This mismatch may restrict the model’s ability to fully acquire and internalize multi-step decision-making patterns, such as delayed reward reasoning, long-horizon credit assignment, and state-dependent strategy adjustments across successive interactions. (2) Offline approximation of real interaction dynamics. Although we introduce process reweighting in the RFT framework to approximate an offline multi-step webpage interaction process, this approach still cannot completely reproduce the complexity, stochasticity, and distribution shift present in real-world environments (e.g., diverse user behaviors, rapidly changing content, and dynamic platform policies). As a result, the learned policy may not generalize optimally to rare or evolving scenarios. Incorporating an online reinforcement learning framework—where the model can interact with the environment and learn from real-time feedback—could better capture these dynamics and more effectively address the above challenges.
@@ 8 Ethics Statement
In this work, we ensure ethical compliance through careful data sourcing and privacy-preserving processing. All data are collected exclusively from publicly accessible websites and used in accordance with applicable laws and platform policies. The resulting dataset contains no personally identifiable information (PII) or other sensitive content. Moreover, we apply strict anonymization and de-identification procedures—such as removing or masking any potentially identifying fields and performing necessary data filtering—to protect user privacy and mitigate potential ethical risks.
We comply with the licensing terms of the public datasets and code repositories used, as specified in their official licenses and terms of service.
Appendix A Future Work
Given the limitations mentioned above, we plan to address them in future work. We note that Shi et al. (2025) proposes a mobile GUI agent framework that leverages reinforcement learning in an online environment. We intend to adapt this framework to web-based GUI agents, enabling the model to learn directly from real-time interactions. Through this approach, GPU memory constraints can be alleviated and the model’s multi-step decision-making capabilities can be further enhanced. Additionally, we plan to build a high-concurrency cluster of browser environments to collect more diverse and complex multi-step instances, further enriching RISK-Data.
Appendix B Task Definition
@@
B.1 External Information Retrieval and Extraction for Risk Intelligence
@@
Table 8: All actions and their definitions used in RISK (Part 2).
Action Definition
@@
B.2 Website Authenticity Verification for Risk Intelligence
Appendix C Data Statistics
C.1 Action Type Distribution
We provide the action type distribution in RISK-Data in Figure 6.
C.2 Multi-step Trajectory Statistics
Appendix D Action Definition
Appendix E Visualization of Weight Curve for Process Reweight
@@ Appendix F Experimental Details
Appendix G Reward Design Analysis
G.1 Process Reweight: Critical Steps in Task Completion.
The goal of process reweighting is to emphasize the importance of later steps in a trajectory, which are more distinguishable and determine the success of task completion. We conduct hyperparameter sensitivity analysis for process reweight, as shown in Figure 9. $\gamma$ controls the weight of the initial step and $\delta$ controls the growth rate of the weight curve, where the visualization is shown in Figure 8. Comparison results reveal that an appropriate setting of process reweight can guide the model to focus on critical steps. However, an excessively low $\gamma$ (e.g., 0.4) or an excessively high $\delta$ (e.g., 7) leads to performance degradation, as the model may neglect the importance of early and intermediate steps, leading to suboptimal learning outcomes.
Table 9: Difficulty measurement analysis.
Measurement Single-step Multi-step
No Reweighting 86.7 79.6
Rule Score 86.1 (-0.6) 78.0 (-1.6)
LLM Response 88.3 (+1.6) 82.8 (+4.8)
G.2 Level Reweight: Difficulty Grading Matters.
Table 10: Difficulty weight configurations.
Configuration Single-step Multi-step
{0.4,0.7,1.0} 87.1 80.9
{0.8,0.9,1.0} 87.8 82.0
{1.0,1.1,1.2} 88.3 82.8
{1.0,1.3,1.5} 88.1 82.4
{1.0,2.0,3.0} 86.6 80.5
In RISK-R1, we use the advanced MLLM’s accuracy to grade the difficulty of samples and set the level reweight accordingly. To validate the effectiveness of this grading method, we compare it with no reweighting and rule-based scoring methods, as shown in Table 9. The rule-based scoring method assigns difficulty levels based on the tool count at each step, where easy, moderate, and difficult samples contain 1, 2, and more than 2 tools, respectively. The results indicate that inappropriate difficulty grading methods can negatively impact model performance, with the rule-based scoring method leading to decreases of 0.6% and 1.6% in single-step and multi-step tasks, respectively. Another notable setting is the difficulty weight configuration, which influences the emphasis on challenging samples. As shown in Table 10, the configuration {1.0, 1.1, 1.2} yields the best performance, while both overly flat ({0.8, 0.9, 1.0}, {0.4, 0.7, 1.0}) and overly steep ({1.0, 1.3, 1.5}, {1.0, 2.0, 3.0}) configurations lead to performance drops. This analysis highlights the importance of appropriate difficulty grading and weight configuration in guiding the reinforcement learning optimization.
Appendix H Ablation Study
@@
Appendix I Cost and Time Analysis
In industrial risk management, current solutions primarily rely on SOTA commercial model APIs integrated into agentic workflows. However, transitioning from proprietary APIs to localized small-scale models is driven by two critical factors: Operational Efficiency (Inference Latency) and Financial Sustainability (Monetary Cost).
Operational Efficiency: High-throughput business tasks are highly sensitive to latency. When utilizing SOTA commercial APIs for a typical 10-step reasoning task, the end-to-end execution time often exceeds 200 seconds due to network overhead and model inference complexity. Such high latency is impractical for real-time business applications where user experience is paramount. By deploying RISK-R1-7B, we significantly reduce the execution time to under 60 seconds, representing a 3x improvement in efficiency that meets the rigorous demands of industrial production.
Financial Sustainability: The cost of SOTA commercial models is becoming increasingly prohibitive for large-scale deployment. For instance, Claude Opus 4.5 Anthropic (2024) involve substantial token-based pricing ( 25/MTok for output tokens). In risk management scenarios requiring millions of daily invocations, these costs are not sustainable. Our research demonstrates that by using the proposed training regimen, a small-scale model (7B) can serve as a cost-effective and high-performance alternative to expensive commercial APIs, maintaining high accuracy while drastically reducing operational expenditure.
Appendix J Error Analysis
We have categorized the failed runs into three primary types to understand the current technical bottlenecks:
Tool/Parameter Errors (Instruction Following): The agent occasionally provides incorrect parameter formats or invokes the wrong tool. This is primarily attributed to the limitations in the small-scale size LLM’s instruction-following capabilities, especially when dealing with complex, multi-layered tool specifications.
Contextual Hallucination (Context Length Constraints): In certain extended trajectories, the agent "hallucinates" or misremembers the results of previous steps. This is largely due to the inherent context window limitations of the small-scale size LLMs, which makes it difficult to maintain perfect coherence across very long tool-calling chains.
Generalization to Unseen Page States: The agent encounters novel UI layouts or system states that were not present in the training/fine-tuning data. This "out-of-distribution" issue remains a significant challenge for robust automation in dynamic business environments.